OpenStudy (anonymous):

please help me in this very hard problem? 1. x-4/x^2-2x-8 - x^2/x^2-16 2. 9/28 + 3/x+2= 3/4

OpenStudy (saifoo.khan):

Factor.

OpenStudy (anonymous):

What are you asked for?. What are the directions of the problem?

OpenStudy (anonymous):

yes is hard, i dont even know, if you guys can help me

OpenStudy (anonymous):

well i dont even know how to start i have done some, But this one i dont even know

OpenStudy (saifoo.khan):

x-4/x^2-2x-8 - x^2/x^2-16 \[\frac{x-4}{x^2 - 2x - 8} - \frac{x^2}{x^2 - 16}\] \[\frac{x-4}{(x-4)(x+2)} - \frac{x^2}{(x-4)(x+4)}\] can u solve now?

OpenStudy (anonymous):

no, I don't understand. But anyways thanks for the help

OpenStudy (saifoo.khan):

x^2 - 2x - 8 can be written as, (x-4)(x+2)